HUBBLE SPACE TELESCOPE DAILY REPORT #1343 PERIOD COVERED: 0000Z (UTC) 03/08/95 - 0000Z (UTC) 03/09/95 Daily Status Report as of 068/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1 5595 (GHRS ECH-A Wavelength Monitor: Cycle 4) The High Resolution Spectrograph was used to make wavelength spectral lamp calibration observations to measure the instrumental characteristics of the Echelle-A spectrograph. The observations were completed as planned, and no problems were reported. 1.2 Completed Two Sets of WF/PC-2 6140 (WF/PC-2 Cycle 4 Calibration: Ramp filters, Internal Calibration) The WF/PC-2 was used to make Earth calibration images as part of a ramp filter calibration test. The observations were completed as planned and no problems were reported. 1.3 Completed WF/PC-2 5302 (The Size and Morphology of Normal Galaxies at Z ~ 3) The WF/PC-2 was used to make Lyman_Alpha emission observations of galaxy PKS0528-25 which is known to have a redshift of Z ~ 3. It is hoped that the nature of this source can be determined. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed GHRS/1 5276 (GHRS Side 1 Detector Calibration : Cycle 4) The High Resolution Spectrograph (Side One) was used to conduct a standard detector test that will be used as part of the routine calibration for the HRS. The observations were completed as planned, and no anomalies were reported. 1.5 Completed FOS/BL 5444 (A UV Spectral Library of Metal-Poor Massive Stars) The FOS (Blue Side) was used for a snapshot program to obtain ultraviolet spectra of AV378 a metal-poor star in the LMC. It will be useful for stellar atmosphere and stellar evolution studies. The observations completed without incident. 1.6 Completed WF/PC-2 5353 (Low-Mass Binaries in the Hyades: Cycle 4 High) The WF/PC-2 was used to make images of the low-mass, late M-Class dwarf star RHY166 which is located in the Hyades cluster. Some of its companion stars may be brown dwarfs. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Notes: The ESTR #2 diagnostic test was completed at 068/0345Z. Analysis is currently being conducted by the Data Capture Facility. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-